# MMAC-Copilot: Multi-modal Agent Collaboration Operating System Copilot

### v2

@@ 3.2 Communication Protocols
Following established multi-agent frameworks [17, 22, 28, 52, 56], MMAC-Copilot adopts Structured Communication Interfaces, avoiding the traps of unconstrained nature language communication. Each agent communicates through a defined schema appropriate for their role.
In specific, we define the output format of the agents as JSON, and stipulate the keys which should be output according to the different roles of each agent. Meanwhile, each agent only receives the content of the required key. For example, "observation" key is exclusive output by Mentor and Viewer, as it is specifically to their role in emphasizing visual observations of screenshots. Apart from themselves, only the Planner receives the content of "observation" key, both when formulating initial plans and when revising them. These protocols are designed to prevent common issues such as data ambiguity in complex multi-agent interactions. By standardizing the communication format and defining clear protocols for message acknowledgment, MMAC-Copilot minimizes the risks of miscommunication and ensures which tasks are executed reliably. Additionally, the use of JSON allows for easy integration of new agents as the system scales, supporting the framework’s adaptability and expansion.
@@ 3.3 Team Collaboration Chain
Team collaboration chains leverage the strengths of dispersed management structures to facilitate enhanced cooperation among team members across different domains. By distributing responsibilities and decision-making power, these chains enable a more agile and responsive organizational environment. Such chain has been successfully implemented across multiple domains [5, 19, 32, 50].
@@ 3.3.1 Dynamic Planning and Subtask Refinement
During the planning phase, Planner uses input from the Viewer to formulate an initial plan. Initial plan often outline the subtasks in a coarse manner, addressing the dynamic nature of visual environments and the inherent limitations of planning based on textual information alone. This initial planning stage sets the foundation for task execution but requires further refinement to address the specifics of the visual statement.
@@ 3.3.2 Continuous Feedback and Iterative Improvement
If discrepancies or incomplete tasks are identified, Mentor generates insights and structured feedback based on a deep analysis of the outcomes. This feedback is crucial for revising the current plans and is incorporated into the next round of planning. This feedback loop between the Mentor and the Planner not only ensures high-quality task completion but also facilitates adaptive learning, allowing the MMAC-Copilot system to evolve its strategies based on real-world interactions and outcomes.
@@ 4.1 Benchmark and setting
We evaluate the performance of MMAC-Copilot using two distinct benchmarks designed to challenge and assess the capabilities of AI assistants in diverse interaction environments. The first is the General AI Assistant Benchmark (GAIA) [26], which includes 466 rigorous question-answering tasks. GAIA primarily tests AI assistants in a question-answering (QA) format, focusing on their ability to utilize API calls for information retrieval and task execution. However, GAIA restricts the evaluation to textual interactions and API-based actions, which do not fully capture the broader range of GUI interactions that are prevalent in real-world applications.
To address these limitations and to evaluate the agents’ capabilities in a more complex and visually oriented environment, we introduced the Visual Interaction Benchmark (VIBench). VIBench comprises cases from three categories of applications: 3D gaming, recreation, and office environments. Unlike GAIA, VIBench emphasizes the necessity for agents to interact with applications through graphical user interfaces (GUIs) that cannot typically be controlled via standard APIs such as Win32 APIs.
Mathematical Representation of VIBench
The core interaction process in VIBench can be mathematically described as follows: Let $S_{0}$ represent the initial system state which includes the current GUI appearance and available actions. Given a user request $R$, the system must transition through a series of states $S_{1},S_{2},\ldots,S_{n}$ via actions $a_{1},a_{2},\ldots,a_{n}$, ultimately achieving the final state $S_{f}$ that satisfies $R$. This process is represented by the sequence:
@@
VIBench thus complements GAIA by adding a layer of complexity and realism, challenging agents to not only respond correctly in a QA format but also to effectively navigate and interact within software environments that mirror everyday user challenges.Examples are shown in Figure 4
@@
To ensure a fair comparison with previous work, we employed the same versions of the ChatGPT API used in prior studies. Specifically, we utilized GPT-4-turbo-1106 and gpt-4-vision-preview. For the Video Analyst, we used gemini-1.5-pro-vision.
@@ 4.2 Evaluation Method
For our VIBench, we adopt human evaluation which follow the SIMA evaluation method [34], recognizing that in numerous instances, task success cannot be automatically inferred. We engage human judges who possess expert-level familiarity with the specific applications, defined as having at least 20 hours of application usage. These experts are asked to assess recorded videos to determine whether the user requests have been accurately executed. During the assessment, human experts are instructed to disregard any irrelevant actions performed by the agents. Their judgments are solely based on whether the stated objectives of the tasks are achieved, without considering the intermediate steps or unrelated actions. Meanwhile, to minimize human subjectivity, we have imposed a maximum limit of 30 execution rounds for task. If the task is not completed within 30 rounds, it is judged as a failure.
@@ 4.3 Baseline
For GAIA, we present the performance of GPT-3.5 and GPT-4 with and without manually set plugins, along with results from AutoGPT, Chamonmile and FRIDAY which all use GPT-4 as their backend. The selection of Plugins for GPT-4 is conducted manually, with individual choosing based on specific user query. For comparative purposes, we also include data on human performance sourced from GAIA [26].
@@ 4.4 Results
The results of our experiments with MMAC-Copilot provide a comprehensive assessment of its capabilities in handling complex, multi-modal tasks with in digital environments. In this session, we present and discuss the findings from our evaluations on both the GAIA benchmark and our newly developed Visual Interaction Benchmark(VIBench).
Model Level 1 Level 2 Level 3 Average
Human* 93.90 91.80 87.30 91.00
GPT-3.5 [1] 4.3 1.89 2.08 2.67
GPT-4 [1] 9.68 1.89 0.00 4.00
AutoGPT4 [33] 15.05 0.63 0.00 5.00
GPT4 Turbo [1] 9.68 6.92 0.00 6.67
GPT-4 Plugins [1] 30.30 9.70 0.00 14.60
Chamomile [15] 16.13 17.61 2.08 14.67
FRIDAY [44] 40.86 20.13 6.12 24.25
Ours 45.16 20.75 6.12 25.91 (+6.8%)
@@
Table 2: Evaluation results on VIBench.
Model 3D Gaming Recreation Office Average
UFO [15] 0.00 28.57 15.38 14.65
FRIDAY [44] 31.58 42.86 30.77 35.07
Ours 63.16 69.23 78.57 70.32 (+35.25)
In VIBench as shown in Table LABEL:tab:VIBench, MMAC-Copilot outperformed other systems across all application categories (3D gaming, recreation, and office scenarios). Our model achieved average scores fo 63.16% in 3D gaming, 69.23% in recreation, and 78.57% in office applications, with an overall average of 70.32%. It indicates MMAC-Copilot’s robust capability in navigating and interacting within complex GUI environments, where direct API interaction is not possible. MMAC-Copilot’s superior performance can be attributed to its effective utilization of image and video analysis, enabling precise navigation and interaction through the GUIs of different software.
@@ 5 Conclusion
This paper presents MMAC-Copilot, a novel framework designed to enhance the interaction capabilities of autonomous virtual agents with operating systems. Our approach capitalizes of autonomous virtual agents with operating systems. Our approach capitalizes on a team collaboration chain that harnesses the collective expertise of each agents, including "Planner", "Librarian", "Programmer", "Viewer", "Video Analyst", and "Mentor". Each agent brings a unique set of skills and perspectives, enabling an adaptive interaction with complex digital environments. Also MMAC-Copilot innovatively mitigates the issues with hallucination among the agents. By facilitating dynamic and team collaboration among the agents, our framework allows for continuous adaptation and refinement of plans based on environmental changes. The performance of MMAC-Copilot has been tested on GAIA benchmark and our newly introduced VIBench, showing significant improvements over existing state-of-the-art systems.
6 Discussion
We acknowledge that while MMAC-Copilot has achieved a level of generality in application through the integration of GUI models for visual detection, it is still limited by the GUI model’s ability to understand UI components. Consequently, MMAC-Copilot struggles with interpreting complex UI interfaces and performing multi-step operations that depend on such complexity. Additionally, MMAC-Copilot’s capabilities in real-time 3D gaming environments reveal gaps in the spatial understanding, which is cruical for tasks such as navigation, aiming and shooting in 3D environments. The framework’s longer inference time impair the ability to perform tasks with higher temporal sensitivity.

### v3

@@ 3.2 Communication Protocols
Following established multi-agent frameworks Hong et al. (2023a); Li et al. (2023); Park et al. (2023); Zhang et al. (2023a); Zhuge et al. (2023), MMAC-Copilot adopts Structured Communication Interfaces, avoiding the traps of unconstrained nature language communication. Each agent communicates through a defined schema appropriate for their role.
In specific, we define the output format of the agents as JSON, and stipulate the keys which should be output according to the different roles of each agent. Meanwhile, each agent only receives the content of the required key. For example, "observation" key is exclusive output by Mentor and Viewer, as it is specifically to their role in emphasizing visual observations of screenshots. Apart from themselves, only the Planner receives the content of "observation" key, both when revising initial plans. These protocols are designed to prevent common issues such as data ambiguity in complex multi-agent interactions. By standardizing the communication format and defining clear protocols for message acknowledgment, MMAC-Copilot minimizes the risks of miscommunication and ensures which tasks are executed reliably. Additionally, the use of JSON allows for easy integration of new agents as the system scales, supporting the framework’s adaptability and expansion.
@@ 3.3 Team Collaboration Chain
Team collaboration chains leverage the strengths of dispersed management structures to facilitate enhanced cooperation among team members across different domains. By distributing responsibilities and decision-making power, these chains enable a more agile and responsive organizational environment. Such chain has been successfully implemented across multiple domains Bang and Dalsgaard (2005); Huh et al. (2019); Selvanathan and Poravi (2018); Yatsuka et al. (2020).
@@ 3.3.1 Dynamic Planning and Subtask Refinement
During the planning phase,In the planning phase, the planner formulates an initial plan based on user requests. Initial plan often outline the subtasks in a coarse manner, addressing the dynamic nature of visual environments and the inherent limitations of planning based on textual information alone. This initial planning stage sets the foundation for task execution but requires further refinement to address the specifics of the visual statement.
@@ 3.3.2 Continuous Feedback and Iterative Improvement
If discrepancies or incomplete tasks are identified, Mentor generates insights and structured feedback based on a deep analysis of the outcomes. This feedback is crucial for revising the current plans and is incorporated into the next round of planning.
@@ 4.1 Benchmark and setting
Model Level 1 Level 2 Level 3 Average
Human* 93.90 91.80 87.30 91.00
Search engine 7.4 0.00 0.00 2.47
GPT-3.5 Achiam et al. (2023) 4.3 1.89 2.08 2.67
GPT-4 Achiam et al. (2023) 9.68 1.89 0.00 4.00
AutoGPT4 Significant-Gravitas (2023) 15.05 0.63 0.00 5.00
GPT4 Turbo Achiam et al. (2023) 9.68 6.92 0.00 6.67
GPT-4 Plugins Achiam et al. (2023) 30.30 9.70 0.00 14.60
Chamomile Guillot (2023) 16.13 17.61 2.08 14.67
FRIDAY Wu et al. (2024) 40.86 20.13 6.12 24.25
Ours 45.16 20.75 6.12 25.91 (+6.8%)
Table 2: Evaluation results on VIBench.
Model 3D Recreation Office Average
UFO 0.00 28.57 15.38 14.65
FRIDAY 31.58 42.86 30.77 35.07
Ours 63.16 69.23 78.57 70.32 (+35.25)
We evaluate the performance of MMAC-Copilot using two distinct benchmarks designed to challenge and assess the capabilities of autonomous agent in diverse interaction environments. The first is the General AI Assistant Benchmark (GAIA) Mialon et al. (2023), which includes 466 rigorous question-answering tasks. GAIA primarily tests AI assistants in a question-answering (QA) format, focusing on their ability to utilize API calls for information retrieval and task execution. However, GAIA restricts the evaluation to textual interactions and API-based actions, which do not fully capture the broader range of GUI interactions that are prevalent in real-world applications.
To address these limitations and to evaluate the agents’ capabilities in a more complex and visually oriented environment, we introduced the Visual Interaction Benchmark (VIBench). VIBench comprises cases from three categories of applications: 3D gaming, recreation, and office environments. Unlike GAIA, VIBench emphasizes the necessity for agents to interact with applications through graphical user interfaces (GUIs) that cannot typically be controlled via standard APIs such as Win32 APIs. The core interaction process in VIBench can be mathematically described as follows: Let $S_{0}$ represent the initial system state which includes the current GUI appearance and available actions. Given a user request $R$, the system must transition through a series of states $S_{1},S_{2},\ldots,S_{n}$ via actions $a_{1},a_{2},\ldots,a_{n}$, ultimately achieving the final state $S_{f}$ that satisfies $R$. This process is represented by the sequence:
@@
VIBench thus complements GAIA by adding a layer of complexity and realism, challenging agents to not only respond correctly in a QA format but also to effectively navigate and interact within a application environments that mirror everyday user challenges. Examples are shown in Figure 3. We also present the detailed list of request utilized in VIBench in Appendix D.1.
@@
To ensure a fair comparison with previous work, we employed the same versions of the ChatGPT API used in prior studies. Specifically, we utilized GPT-4-turbo-1106 and gpt-4-vision-preview. For the Video Analyst, we used gemini-1.5-pro-vision. More detailed experiment setting could been seen in Appendix C.
@@ 4.2 Evaluation Method
For our VIBench, we adopt human evaluation which follow the SIMA evaluation method SIMA Team (2024), recognizing that in numerous instances, task success cannot be automatically inferred. We engage human judges who possess expert-level familiarity with the specific applications, defined as having at least 20 hours of application usage. These experts are asked to assess recorded videos to determine whether the user requests have been accurately executed. During the assessment, human experts are instructed to disregard any irrelevant actions performed by the agents. Their judgments are solely based on whether the stated objectives of the tasks are achieved, without considering the intermediate steps or unrelated actions. Meanwhile, to minimize human subjectivity, we have imposed a maximum limit of 30 execution rounds for task. If the task is not completed within 30 rounds, it is judged as a failure.
@@ 4.3 Baseline
For GAIA, we present the performance of GPT-3.5 and GPT-4 with and without manually set plugins, along with results from AutoGPT, Chamonmile and FRIDAY which all use GPT-4 as their backend. The selection of Plugins for GPT-4 is conducted manually, with individual choosing based on specific user query. For comparative purposes, we also include data on human performance sourced from GAIA Mialon et al. (2023).
@@ 4.4 Results
In VIBench as shown in Table 2, MMAC-Copilot outperformed other systems across all application categories (3D gaming, recreation, and office scenarios). Our model achieved average scores of 63.16% in 3D gaming, 69.23% in recreation, and 78.57% in office applications, with an overall average of 70.32%. It indicates MMAC-Copilot’s robust capability in navigating and interacting within complex GUI environments, where direct API interaction is not possible.
@@ 5 Conclusion
MMAC-Copilot is a novel framework enhances agent interactions within complex digital environments. This framework effectively addresses agent hallucination, ensuring reliable responses to environmental changes. Tested on the GAIA benchmark and our VIBench, MMAC-Copilot demonstrates superior performance compared to existing systems.
Limitation
We acknowledge that while MMAC-Copilot has achieved a level of generality in application through the integration of GUI models for visual detection, it is still limited by the GUI model’s ability to understand UI components. Consequently, MMAC-Copilot struggles with interpreting complex UI interfaces and performing multi-step operations that depend on such complexity. Additionally, MMAC-Copilot’s capabilities in real-time 3D gaming environments reveal gaps in the spatial understanding, which is crucial for tasks such as navigation, aiming and shooting in 3D environments. The framework’s longer inference time impair the ability to perform tasks with higher temporal sensitivity.
@@
Appendix A Demo on Mac OS
The following Figure 4 illustrates an example of MMAC-Copilot performing an action sequence on macOS in response to the user request “Play <Love Story> on Spotify.” Each subfigure represents a distinct step in the process: (A) shows the “Planner” formulating the initial plan; (B) the “Programmer” launching the application using a bash command; (C) the “Viewer” clicking the search box and entering <Love Story>; and (D) the “Viewer” clicking the play button. This demonstrates how MMAC-Copilot efficiently handles a media playback request from the user.
Appendix B Agent prompts
To better explain the operating mechanism of Programmer, we have provided the complete prompts in the Figure 9,10,11,12,13,15,16,17 and 18.
Appendix C GAIA Detail Experiment Setting
GAIA serves as a comprehensive benchmark for evaluating general AI assistants by testing their ability to handle real-world scenarios across various capabilities, including advanced reasoning, multi-modality handling, coding, and web browsing. The benchmark comprises text-based questions, sometimes accompanied by supplementary files such as images or spreadsheets. These questions are designed to cover a broad range of use cases from everyday tasks to complex scientific queries, necessitating concise and unambiguously correct answers that are straightforward to verify.
C.1 Levels of Difficulty
The questions in GAIA are categorized into three levels based on their complexity:
  • Level 1: Simplest tasks, generally requiring minimal tool use and up to five steps.
  • Level 2: More complex tasks that require integrating information from multiple sources, involving about 5 to 10 steps.
  • Level 3: The most complex tasks, demanding extensive sequences of actions and the use of multiple tools to solve real-world problems.
We present a sample GAIA question from Level 1 through Level 3 in Figure 5
Figure 5: Sample Questions from GAIA. These tasks necessitate core competencies including reasoning, the ability to handle multiple modalities, and proficiency in using tools. The answers to these questions are unequivocally defined and are intentionally unlikely to be present in unstructured training datasets. Several questions are supplemented with visual evidence, mirroring authentic scenarios and enhancing the precision of query formulation.
C.2 Evaluation Methodology
GAIA employs an automated, precise, and factual evaluation method. Each question demands an answer in a specified format (string, number, or comma-separated list) and correctness is verified through an exact match with a pre-defined ground truth, ensuring the evaluation’s reliability and efficiency.
C.3 Model Experimental Setup
In this section, we introduce the detailed experiment setting of our comparison models.
We evaluates leading AI models, including GPT-3.5, GPT4 and AutoGPT, also conduct non-AI baselines experiments such as human annotators and traditional web search methods. We will introduce each experiment setting as following.
Comparison Models:
GPT3.5/GPT4:
This setup evaluates the core linguistic and reasoning capabilities of GPT4 without the aid of any external tools or plugins. We use a prefix prompt before asking the model a question. To ease answer extraction, we specify a format in the prefix prompt which could been saw in Figure 6.
GPT4 with Plugins:
In this configuration, GPT4 is enhanced with selected plugins that extend its functionality. This includes tools for advanced data analysis, web browsing, and handling of third-party APIs.
AutoGPT:
Utilizing GPT4 as its backend, AutoGPT automatically selects plugins based on the context of the query. This model is tested for its efficiency in tool selection and its ability to autonomously adapt to complex query requirements.
Human*:
Consisting of experts in various fields, human annotators provide responses based on their expertise and research abilities. Their performance sets a high standard for AI models, especially in tasks requiring deep domain knowledge and nuanced understanding.
Search engine:
Using conventional search engines, this baseline evaluates the effectiveness of straightforward information retrieval. Specifically, we enter our questions into a search engine and review the first page of results to see if the answer can be inferred.
Notes: Each model undergoes rigorous testing across a series of structured tasks that range from simple to highly complex, categorized into three levels of difficulty as previously described. Models are evaluated based on accuracy, the relevance of the responses, and the efficiency of handling the queries. Each session is replicated three times to ensure consistency and reliability in the results.
Figure 6: The system prompt to answer GAIA.
Appendix D VIBench
D.1 VIBench Results List
In Table 3, we provide a comprehensive list of the user requests featured in VIBench, along with the corresponding performance of MMAC-Copilot. These requests cover a range of commonly used functions across eight popular applications. For clarity, requests that involve follow-up tasks are numbered sequentially. In the Success column, “” denotes that MMAC-Copilot completes the request successfully, while “” indicates a failure to fufill the request.
Table 3: The requests in VIBench and detailed results achieved by MMAC-Copilot
Operation Application Success
Ride on vehicle [Press F] PUBG
Probe shot [Press Left mouse] PUBG
Picking up supplies [Identify the screen and press F] PUBG
Assemble accessories and discard unnecessary ones PUBG
Use E skill Valorant
Shoot Valorant
Climb up the rope Valorant
Absorb energy ball Valorant
Take turns to release Q skills Genshin
Pick up the Divine Eye Genshin
Freeze trap Genshin
Fishing Genshin
Climb Genshin
Send message to xxx DingTalk
Create meeting DingTalk
Create a group DingTalk
End meeting TencentMeeting
Turn on camera TencentMeeting
Create a meeting TencentMeeting
Join a meeting TencentMeeting
Click the record bottom TencentMeeting
Copy Invitation TencentMeeting
Open the library in steam Steam
Open the friends interface Steam
Click on store and then click on the search box Steam
View current music and play it Spotify
Search for a specific piece of music and save it Spotify
Open docuseries in Netflix Netflix
Save XXX Series to my list Netflix
Remove XXX from my list Netflix
D.2 Visualization
We have visualized VIBench with requests in Figure 7 and 8.
D.3 Adapt for Windows
To compatible with the Windows platform, we made some adaptations focusing on tool integration and modifications for platform-specific libraries.
One of the primary adaptations involved the integration of tools such as the Python environment and Windows Shell support. The Python interpreter was configured to work within the Windows platform and Windows path handling were incorporated to ensure smooth execution of scripts and Python modules. In addition to tool integration, some Python libraries required adaptation to function properly on Windows. We used alternative Python modules to replace database libraries that did not fully support the Windows environment for tool persistent. As Python libraries behave differently on Windows due to variations in concurrent requests management, we also made some adaptation for it.
Figure 7: How VIBench works(Part-I)
Figure 8: How VIBench works(Part-II)
Figure 9: Prompt for Programmer Part
Figure 10: Prompt for Programmer Part
Figure 11: Prompt for Programmer Part
Figure 12: Prompt for Programmer Part
Figure 13: Prompt for Programmer Part
Figure 14: Prompt for Programmer Part
Figure 15: Prompt for Programmer Part
Figure 16: Prompt for Programmer Part
Figure 17: Prompt for Programmer Part
Figure 18: Prompt for Programmer Part